# A spectral method for multi-view subspace learning using the product of projections

### v1

@@ B.1. Simulations on synthetic data for more than two views
    Table 4. Quality of the joint and individual subspace estimates in terms of the metric (6), where larger values indicate better performance. All values are multiplied by $10$. The results are averaged across 100 iterations. Maximum standard error of the mean is $0.03$. We highlight the best result in bold and underline second best result. We denote by () methods that do not allow supplying marginal ranks for each view. We denote Lock et al. (2013) as JIV, Feng et al. (2018) as AJI, Gaynanova & Li (2019) as SLI, Park & Lock (2020) as UNI, and our method as PPD.
@@ B.2. Ablation study for bootstrap
    We modify the bootstrap method proposed in Prothero et al. (2024) to better align the resampled singular vector bases with the observed data. Specifically, given orthogonal resampled bases $U_{1,(b)}$ and $U_{2,(b)}$, we adjust $U_{2,(b)}$ to align with $U_{1,(b)}$ using the formula:$U_{2,(b)}\leftarrow U_{1,(b)}\cos(\Sigma_{\widehat{M}})+U_{2,(b)}\sin(\Sigma_{\widehat{M}}),$ where $\Sigma_{\widehat{M}}$ is a diagonal matrix containing the principal angles between the column spaces $\operatorname{col}(\widehat{X}_{1})$ and $\operatorname{col}(\widehat{X}_{2})$. This alignment ensures that the resampled bases reflect the geometric relationship observed in the data, which is crucial for accurate estimation of $\varepsilon_{1}$ and $\varepsilon_{2}$. Without this correction, a straightforward approach would be to randomly sample $U_{1,(b)}$ and $U_{2,(b)}$ from the Haar measure. However, as we argue in Section 3.3, such an approach tends to underestimate $\varepsilon_{1}$ and $\varepsilon_{2}$, particularly when the rank-to-dimension ratio is low. In this regime, randomly sampled bases are likely to be nearly orthogonal, which does not capture the true alignment between the column spaces of $\widehat{X}_{1}$ and $\widehat{X}_{2}$.
@@ B.3. Empirical error of the bootstrap
    We simulate the data following Section 4, setting. We run our bootstrap model for a range of signal-to-noise ratios and angles, supplying the model with the true marginal ranks. We report the difference $\hat{\epsilon}_{1}-\epsilon_{1}$ in Figure 6. From the figure, we see that our bootstrap estimator tends to be accurate, except for the low signal-to-noise ratios for which it overestimates the true parameter. We investigate this issue and find it related to the overestimation of the noise variance by the robust estimator $y_{\text{med}}/\mu_{\beta}$ with $y_{\text{med}},\mu_{\beta}$ denoting the median singular value of $Y$ and the median of the Marchenko-Pastur distribution with parameter $\beta$. We show the error in the variance estimation in Figure 6. Recall that in our bootstrap as described in Section 3.3, we impute the signal directions back into the noise estimate $\check{E}_{k}$, which are scaled by the noise variance. As our estimate of the noise variance overestimates the truth, correspondingly our noise estimate becomes inflated, which finally leads to the inflated $\widehat{\varepsilon}_{1}$ estimate. We leave the improvement of our bootstrap estimator in the low signal-to-noise ratio regime as future work.

### v2

@@ B.1. Simulations on synthetic data for more than two views
    Table 4. Quality of the joint and individual subspace estimates in terms of the metric (6), where larger values indicate better performance.
@@ B.2. Ablation study for bootstrap
[…]
@@ B.3. Empirical error of the bootstrap
    We simulate the data following Section 5, setting. We run our bootstrap model for a range of signal-to-noise ratios and angles, supplying the model with the true marginal ranks. We report the difference $\hat{\epsilon}_{1}-\epsilon_{1}$ in Figure 6. From the figure, we see that our bootstrap estimator tends to be accurate, except for the low signal-to-noise ratios for which it overestimates the true parameter. We investigate this issue and find it related to the overestimation of the noise variance by the robust estimator $y_{\text{med}}/\mu_{\beta}$ with $y_{\text{med}},\mu_{\beta}$ denoting the median singular value of $Y$ and the median of the Marchenko-Pastur distribution with parameter $\beta$. We show the error in the variance estimation in Figure 6. Recall that in our bootstrap as described in Section 4, we impute the signal directions back into the noise estimate $\check{E}_{k}$, which are scaled by the noise variance. As our estimate of the noise variance overestimates the truth, correspondingly our noise estimate becomes inflated, which finally leads to the inflated $\widehat{\varepsilon}_{1}$ estimate. We leave the improvement of our bootstrap estimator in the low signal-to-noise ratio regime as future work.